the couple has a son and a daughter. the age of the father is threee times that of his daughters age the age of son is half of his mother. the wife is 9 years younger to her husband and the brother is 7 years older then than his sister. what is the age of the mother?

Answer by mananth(16946)

let mother's age =x
husband = x+9
..
son = x/2
sister 1/3 (x+9)
brother = 1/3(x+9)+7
..
brother = son
..
x/2 = 1/3(x+9)+7
3x/2 = x+9+7
3x/2 = x+16
3x= 2(x+16)
3x=2x+32
3x-2x=32
x= 32 mother's age.